In the laboratory, hydrogen gas is usually made by the following reaction:
Zn(s) + 2 HCl(aq) → H2(g) + ZnCl2(aq)
How many liters of H2 gas, collected over water at an atmospheric pressure of 752 mm Hg and a temperature of 21.0°C, can be made from 1.566 g of Zn and excess HCl? The partial pressure of water vapor is 18.65 mm Hg at 21.0°C.
Answer #1

The moles of H2 produced are calculated:

n H2 = 1.566 g Zn * (1 mol / 65.38 g) * (1 mol H2 / 1 mol Zn) = 0.024 mol

The pressure of H2 is calculated:

P H2 = Total P - P H2O = 752 - 18.65 = 733.35 mmHg

The pressure in atm is calculated:

P H2 = 733.35 mmHg * (1 atm / 760 mmHg) = 0.964 atm

The volume of H2 collected is calculated:

V H2 = n * R * T / P = 0.024 mol * 0.082 atm * L / mol * K * 294 K / 0.964 atm = 0.6 L

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.